OpenStudy (anonymous):

What is the circumference of a circle with an area of 49pi in2

OpenStudy (amistre64):

2pi sqrt(49) i think

OpenStudy (amistre64):

49pi = pi r^2 49 = r^2 sqrt(49) = r .... yep

OpenStudy (anonymous):

24.5 pi inches 14 pi inches 7 pi inches 3.5 piinches

OpenStudy (anonymous):

those r my choices